# Statistical Learning for Heterogeneous Treatment Effects: Pretraining, Prognosis, and Prediction

### v1

@@ Abstract
Robust estimation of heterogeneous treatment effects is a fundamental challenge for optimal decision-making in domains ranging from personalized medicine to educational policy. In recent years, predictive machine learning has emerged as a valuable toolbox for causal estimation, enabling more flexible effect estimation. However, accurately estimating conditional average treatment effects (CATE) remains a major challenge, particularly in the presence of many covariates. In this article, we propose pretraining strategies that leverages a phenomenon in real-world applications: factors that are prognostic of the outcome are frequently also predictive of treatment effect heterogeneity. In medicine, for example, components of the same biological signaling pathways frequently influence both baseline risk and treatment response. Specifically, we demonstrate our approach within the $R$-learner framework, which estimates the CATE by solving individual prediction problems based on a residualized loss. We use this structure to incorporate “side information” and develop models that can exploit synergies between risk prediction and causal effect estimation. In settings where these synergies are present, this cross-task learning enables more accurate signal detection: yields lower estimation error, reduced false discovery rates, and higher power for detecting heterogeneity.
@@ 1 Introduction
For many applications, from personalized medicine to educational policies, understanding treatment effects beyond global averages is critical for optimal decision-making and treatment prioritization. For example, a cancer treatment may show average benefits for a patient population, yet individual responses can vary widely among patients (Tomasik et al., 2024). In the presence of such heterogeneity, the average may be composed of individuals with large, moderate, and negative treatment effects (Kravitz et al., 2004). Those with minimal benefit or harm are better off receiving no or alternative treatments. This explains why in medicine and other high-stakes settings, relying solely on average treatment effects (ATE) can lead to suboptimal decisions–potentially harming some individuals or, incorrectly, withholding treatment from those who could benefit. To advance precision medicine, researchers and practitioners are, therefore, interested in developing as highly personalized treatment recommendations as possible (Kosorok and Laber, 2019). One common strategy for deriving personalized treatment policies is to use a risk-based approach, where individuals prone to poor outcomes receive treatment first (Kent et al., 2016, 2020). An alternative approach is to prioritize individuals based on their treatment benefit (Manski, 2004; Wager and Athey, 2018).
@@
With the widespread availability of large-scale observational data, researchers now have access to rich data environments that offer the potential to estimate the conditional average treatment effects for increasingly granular subgroups. Simultaneously, machine learning has emerged as a useful toolbox, offering several advantages for CATE estimation (Athey and Imbens, 2016). A large body of work has made considerable progress towards more flexible estimation, e.g. by using the lasso (Imai and Ratkovic, 2013), boosting (Powers et al., 2018), multivariate adaptive regression (Powers et al., 2018), neural networks (Shalit et al., 2017), and random forests (Wager and Athey, 2018), to name just a handful of approaches. Moreover, these models can be flexibly integrated into metalearners (Kennedy, 2023; Künzel et al., 2019; Morzywolek et al., 2023; Nie and Wager, 2021; van der Laan et al., 2024) that decompose the exercise of effect estimation into separate prediction tasks. This has enhanced the role of statistical learning for CATE estimation in recent years.
@@
Beyond more flexible estimation, automated CATE models offer a principled and objective way of identifying and reporting subgroups that drive heterogeneous treatment responses (Athey et al., 2019). This provides a remedy to a more traditional approaches in which practitioners need to define relevant subgroups a priori. Moreover, it mitigates practices similar to $p$-hacking, where possible subgroups are repeatedly searched until a putative signal for heterogeneity is found. Especially in areas like genetics, where treatment heterogeneity is often the results of complex interactions, such manual approaches are not scalable and become easily impractical.
@@
Despite these advances, there remains a fundamental statistical difficulty in CATE estimation (Kennedy et al., 2024). Identifying heterogeneous treatment effects is especially challenging when sample sizes are limited (e.g., in rare diseases) or when effects are diffuse and involve complex interactions. Identifying effect modifiers then often resembles the task of finding a “needle in a haystack”. For example, in settings with $p=10$ binary covariates, there are already $10242^{10}=$ 1024 possible subgroups that could explain heterogeneity. The challenge grows in high-dimensional settings like gene expression data, where $p$ can be as large as, or even exceed $n$. In those settings, low power, overfitting and high variance are concerns that complicate estimating the CATE (Hastie et al., 2009; Powers et al., 2018; Candes et al., 2018).
@@
Low sample sizes and high-dimensional data are common challenges in medical settings. For instance, clinical trials increasingly include more covariates (e.g., from molecular data), while their sample sizes have not kept pace. As a result, there is an urgent need for methods that improve upon traditional approaches to CATE estimation. Ignatiadis and Huber (2021) suggest mitigating these challenges by leveraging domain knowledge and hidden “side information” in the data—for example, to derive data-driven weights for hypothesis testing. In a similar vein, Fortney et al. (2015) and Dobriban et al. (2015) show how prior information on disease-related gene loci may help elucidate loci for longevity due to their existing overlap. Seemingly unrelated disease loci can predict treatment responses across multiple conditions. These examples highlight that real-world datasets in biomedicine, though often characterized by a low signal-to-noise ratio, are rich in connections and structures. These might involve risk measures that serve as surrogates for treatment effects and help uncover effect modifiers. In predictive settings, statistical learning methods that incorporate weights from related tasks—commonly known as transfer learning—have gained significant attention over the past decades (Argyriou et al., 2006; Craig et al., 2024; Gu et al., 2024; Obozinski et al., 2010; Zou, 2006; Li et al., 2022). Harnessing putative “side information” for heterogeneous treatment effects, bears a considerable potential, but requires statistical learning strategies that can synergize seemingly distinct prediction tasks for CATE estimation.
@@
In this article, we use the term “pretraining” as an umbrella term to denote various statistical learning procedures that enable a sequential transfer of information from one (or multiple) prediction model(s) to another. Depending on the architecture of each model, this can take the form of using the active set from a lasso to derive penalty factors for the second model (adaptive lasso), performing weighted feature sampling based on feature importance in tree-based learners or transferring weights to a neural network.
@@ 2 Leveraging Side Information for Heterogeneous Treatment Effect Estimation
The ideal clinical decision making quantity is the $i$-th unit’s treatment effect, $Y_{i}(1)-Y_{i}(0)$. By the fundamental problem of causal inference, this quantity, however, is not point identified because each unit $i$ can either receive treatment or not. In order to gain insight into personalized treatment responses it is helpful to instead shift attention to conditional average treatment effects (CATE), defined as $\tau(x)=\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid X_{i}=x\right]$. We denote the mean conditional potential outcome in arm $w$ by $\mu_{(w)}(x)=\mathbb{E}\left[Y_{i}(w)\mid X_{i}=x\right]$. Given Assumption 1, $\mu_{(w)}(x)=\mathbb{E}\left[Y_{i}\mid X_{i}=x,W_{i}=w\right]$, and the CATE is identified by
@@
A drawback of this simple direct approach is regularization bias due to, for example, treatment imbalance (Künzel et al., 2019). Recent advances have produced an array of improvements to this approach by decomposing prediction tasks into tailored loss functions that directly target the CATE (e.g., Foster and Syrgkanis (2023); Kennedy (2023); Künzel et al. (2019); Morzywolek et al. (2023); Nie and Wager (2021); van der Laan et al. (2024)). The goal of these approaches is to strip out confounding and baseline effects and then focus on the CATE target parameter. Common in these loss-based approaches are nuisance components that we might leverage for the final CATE goal. In particular, the $R$-learner (Nie and Wager, 2021) leverages Robinson’s transformation (Robins, 2004; Robinson, 1988) to express the CATE function as
@@
One particularly useful ingredient in this loss is the model for the conditional mean in the absence of treatment, $\hat{\mu}_{(0)}(X_{i})$. If features that are predictive of baseline outcomes also matter for treatment benefits, then sharing information between $\hat{\mu}_{(0)}(X_{i})$ and $\hat{\tau}(X_{i})$ might be advantageous in estimating the CATE. For statistical reasons that we discuss shortly and brevity, we restrict our study to exploring leveraging the $R$-learner to share information between marginalized outcomes $m(X_{i})$ and the CATE. However, note that the same logic holds for the $DR$-learner. For a general discussion and comparison of these orthogonal losses for CATE estimation, we refer to Morzywolek et al. (2023).
@@ 2.1 Shared Support between Prognostic and Predictive Factors
A common observation in medicine and other fields is that factors that are prognostic of an outcome are also predictive of treatment response, i.e. there is a overlap in the support of features that are relevant to solve the objective function for the mean outcome function $m(X_{i})$ and those for the CATE function $\tau(X_{i})$. Equation 6 expresses this phenomenon in an easy toy based on a linear model. The covariate $X_{i1}$ is simply prognostic of the outcome $Y_{i}$ (in the absence of the treatment); $X_{i3}$ is predictive as it interacts with the treatment. Finally, $X_{i2}$ is both prognostic and predictive.
@@
(Prognostic Factor). A prognostic factor is any covariate or combination of covariates that contains information about the likelihood an outcome, independent of any treatment assignment (Ballman, 2015; National Cancer Institute, 2025b).
@@
Figure 1: Overview of the Pretrained $R$-learner. The figure shows the step-wise procedure of the $R$-learner with the mean outcome function and the CATE. The upper panel shows different pretraining strategies. The lower panels illustrate how these pretraining approaches are integrated into the $R$-learner using the mean outcome function as example. The active set for the mean outcome model (left panel) is used for pretraining to inform the estimation of the conditional average treatment effect (CATE) model. Here, gender, weight, and biomarker II are not in the support of the mean outcome model (middle panel) and get penalized at $1/\alpha$ when estimating the CATE model (right panel). Optionally, previously known, expert-validated predictive factors are exempt from any penalties. For example, biomarker II is a known to be predictive (and thus protected from any penalties a priori), but not in the active set of the mean outcome model. Age, tumor size, and biomarker I are in the supports of both models, and do not get penalized through pretraining. Icons from Servier Medical Art, licensed under CC BY 4.0, https://creativecommons.org/licenses/by/4.0/ .
@@
The overlap of predictive and prognostic factor also plays a role in psychiatry and cardiology: patients with depression have a higher risk and poorer outcomes for cardio-vascular diseases (CVD) (Lichtman et al., 2014). Furthermore, patients who suffer from depression are at higher risk for medication non-adherence, which can decrease the effectiveness of CVD medications such as statins and anti-hypertensive treatments (DiMatteo et al., 2000; Carney and Freedland, 2017). Finally, the phenomenon of shared support is likely to extend to many other disciplines outside medicine. Athey et al. (2025) demonstrate that baseline responses are easier to estimate than treatment effect and leverage this for enhanced causal predictions to target behavioral nudges. In education, for example, school-specific characteristics, such as learning environments, not only predict students’ overall educational performance, but also often modify the effects of educational interventions (Yeager et al., 2019).
@@ 2.2 Side Information from Selection Factors
It is possible to extend our concept of pretraining to the propensity score function in observational settings where treatment assignment is the function of some covariates. Suppose clinicians tend to recommend a treatment to patients in an age group where they believe it is most effective. These patients will have high treatment propensity. If clinicians are generally correct, then the same covariates driving propensity will also predict treatment benefit. Cohen and DeRubeis (2018) discuss the role of prognostic, predictive and selection factors in the context of depression. Selection factors also play an important role in phase-II and phase-III cancer trials, where investigators choose one (“physician’s choice”) among many treatment options in the control arm (Modi et al., 2022; Ribas et al., 2015).
@@
(Selection Factor). A selection factor is any pre-exposure covariate which contains information about the likelihood of treatment assignment.
Pretraining on selection factors might be also relevant in other areas for research such as education. There is large body of research examining the effect of learning interventions or college admission on outcomes (e.g. future income). In this example, selection factors such as socio-economic status, motivation, belief’s about one’s future, and health might be strong effect modifiers (Yeager et al., 2016).
@@ 3 Using the Lasso to Leverage Prognostic Side Information
In order to instantiate an $R$-learner that leverages shared support, we need learners for the mean conditional outcome function and CATE function amendable to this support-sharing task. One tractable solution is to leverage an adaptive lasso for the CATE function where the differential shrinkage parameters are set through the lasso active set solution for the mean conditional outcome function. For exposition it is helpful to first consider regularized high-dimensional linear models. Suitable featurizations of the covariate matrix (Hastie et al., 2009) allow us to extend this approach to non-linear settings as revisited in section 4. We start out by assuming a randomized controlled trial setting where the propensity scores are known and omit this estimation task. We start by fitting $m(X_{i})$ with the lasso
@@
where $\lambda_{1}$ is a regularization parameter that minimizes the cross-validated MSE. In addition to the estimated model coefficients $\hat{\beta}$, the lasso yields an active set of features with nonzero coefficients. If some of these coefficients are not part of the active set, then as argued in Section 2.1, they might be less useful for estimating the CATE. In the next step we incorporate information from this active set by estimating $\tau(X_{i})$ via an adaptive lasso using (3)
@@
We can employ a similar strategy in the $DR$-learner (5) using its nuisance parameters for pretraining. While pretraining based on the conditional outcome model for the control arm $\mu_{(0)}(X_{i})$ is appealing, it has one disadvantage: depending on the fraction of treated units, we will have less data at our disposition. Moreover, pretraining using $m(X_{i})$ has the potential to pick up signal that stems from $\tau(X_{i})$ given that under Assumption 1 we have that $m(X_{i})=\mu_{(0)}(X_{i})+e(X_{i})\tau(X_{i})$. This provides a potential safeguard against penalizing predictive factors that are not prognostic when there is little to no overlap between the two types of covariates.

### v2

@@ Abstract
Robust estimation of heterogeneous treatment effects is a fundamental challenge for optimal decision-making in domains ranging from personalized medicine to educational policy. In recent years, predictive machine learning has emerged as a valuable toolbox for causal estimation, enabling more flexible effect estimation. However, accurately estimating conditional average treatment effects (CATE) remains a major challenge, particularly in the presence of many covariates. In this article, we propose pretraining strategies that leverage a phenomenon in real-world applications: factors that are prognostic of the outcome are frequently also predictive of treatment effect heterogeneity. In medicine, for example, components of the same biological signaling pathways frequently influence both baseline risk and treatment response. Specifically, we demonstrate our approach within the $R$-learner framework, which estimates the CATE by solving individual prediction problems based on a residualized loss. We use this structure to incorporate side information and develop models that can exploit synergies between risk prediction and causal effect estimation. In settings where these synergies are present, this cross-task learning enables more accurate signal detection, yields lower estimation error, reduced false discovery rates, and higher power for detecting heterogeneity.
Keywords: Statistical learning, causal inference, heterogeneous treatment effects, biomarkers
@@ 1 Introduction
For many applications, from personalized medicine to educational policies, understanding treatment effects beyond global averages is critical for optimal decision-making and treatment prioritization. For example, a cancer treatment may show average benefits for a patient population, yet individual responses can vary widely among patients (Tomasik et al., 2024). In the presence of such heterogeneity, the average may be composed of individuals with large, moderate, and negative treatment effects (Kravitz et al., 2004). Those with minimal benefit or harm are better off receiving no or alternative treatments. This explains why in medicine and other high-stakes settings, relying solely on average treatment effects (ATE) can lead to suboptimal decisions–potentially harming some individuals or, incorrectly, withholding treatment from those who could benefit. To advance precision medicine, researchers and practitioners are, therefore, interested in developing as highly personalized treatment recommendations as possible (Kessler and Luedtke, 2021; Kosorok and Laber, 2019). One common strategy for deriving personalized treatment policies is to use a risk-based approach, where individuals prone to poor outcomes receive treatment first (Kent et al., 2016, 2020). An alternative approach is to prioritize individuals based on their treatment benefit (Manski, 2004; Wager and Athey, 2018).
@@
With the widespread availability of large-scale observational data, researchers now have access to rich data environments that offer the potential to estimate the conditional average treatment effects for increasingly granular subgroups. Simultaneously, machine learning has emerged as a useful toolbox, offering several advantages for CATE estimation (Athey and Imbens, 2016). A large body of work has made considerable progress towards more flexible estimation, e.g., by using the lasso (Imai and Ratkovic, 2013), boosting (Powers et al., 2018), multivariate adaptive regression (Powers et al., 2018), neural networks (Shalit et al., 2017), and random forests (Wager and Athey, 2018), to name just a handful of approaches. Moreover, these models can be flexibly integrated into metalearners (Kennedy, 2023; Künzel et al., 2019; Morzywolek et al., 2023; Nie and Wager, 2021; van der Laan et al., 2024) that decompose the exercise of effect estimation into separate prediction tasks. This has enhanced the role of statistical learning for CATE estimation in recent years.
@@
Beyond more flexible estimation, automated CATE models offer a principled and objective way of identifying and reporting subgroups that drive heterogeneous treatment responses (Athey et al., 2019). This provides a remedy to more traditional approaches in which practitioners need to define relevant subgroups a priori. Moreover, it mitigates practices similar to $p$-hacking, where possible subgroups are repeatedly searched until a putative signal for heterogeneity is found. Especially in areas like genetics, where treatment heterogeneity is often the result of complex interactions, such manual approaches are not scalable and become easily impractical.
@@
Despite these advances, there remains a fundamental statistical difficulty in CATE estimation (Kennedy et al., 2024). Identifying heterogeneous treatment effects is especially challenging when sample sizes are limited (e.g., in rare diseases) or when effects are diffuse and involve complex interactions. Identifying effect modifiers then often resembles the task of finding a “needle in a haystack”. For example, in settings with $p=10$ binary covariates, there are already $10242^{10}=$ 1024 possible subgroups that could explain heterogeneity. The challenge grows in high-dimensional settings like gene expression data, where $p$ can be as large as, or even exceed $n$. In those settings, low power, overfitting, and high variance are concerns that complicate estimating the CATE (Hastie et al., 2009; Powers et al., 2018; Candes et al., 2018).
@@
Low sample sizes and high-dimensional data are common challenges in medical settings. For instance, clinical trials increasingly include more covariates (e.g., from molecular data), while their sample sizes have not kept pace. As a result, there is an urgent need for methods that improve upon traditional approaches to CATE estimation. Ignatiadis and Huber (2021) suggests mitigating these challenges by leveraging domain knowledge and hidden side information in the data—for example, to derive data-driven weights for hypothesis testing. In a similar vein, Fortney et al. (2015) and Dobriban et al. (2015) show how prior information on disease-related gene loci may help elucidate loci for longevity due to their existing overlap. Seemingly unrelated disease loci can predict treatment responses across multiple conditions. These examples highlight that real-world datasets in biomedicine, though often characterized by a low signal-to-noise ratio, are rich in connections and structures. These might involve risk measures that serve as surrogates for treatment effects and help uncover effect modifiers. In predictive settings, statistical learning methods that incorporate weights from related tasks—commonly known as transfer learning—have gained significant attention over the past decades (Argyriou et al., 2006; Craig et al., 2024; Gu et al., 2024; Obozinski et al., 2010; Zou, 2006; Li et al., 2022). Harnessing putative side information for heterogeneous treatment effects bears considerable potential, but requires statistical learning strategies that can synergize seemingly distinct prediction tasks for CATE estimation.
@@
In this article, we use the term “pretraining” as an umbrella term to denote various statistical learning procedures that enable a sequential transfer of information from one (or multiple) prediction model(s) to another. Depending on the architecture of each model, this can take the form of using the active set from a lasso to derive penalty factors for the second model (adaptive lasso), performing weighted feature sampling based on feature importance in tree-based learners, or transferring weights to a neural network.
@@ 2 Leveraging Side Information for Heterogeneous Treatment Effect Estimation
The ideal clinical decision making quantity is the $i$-th unit’s treatment effect, $Y_{i}(1)-Y_{i}(0)$. By the fundamental problem of causal inference, this quantity, however, is not point-identified because each unit $i$ can either receive treatment or not. In order to gain insight into personalized treatment responses, it is helpful to instead shift attention to conditional average treatment effects (CATE), defined as $\tau(x)=\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid X_{i}=x\right]$. We denote the mean conditional potential outcome in arm $w$ by $\mu_{(w)}(x)=\mathbb{E}\left[Y_{i}(w)\mid X_{i}=x\right]$. Given Assumption 1-2, $\mu_{(w)}(x)=\mathbb{E}\left[Y_{i}\mid X_{i}=x,W_{i}=w\right]$, and the CATE is identified by
@@
A drawback of this simple approach is regularization bias due to, for example, treatment imbalance (Künzel et al., 2019). Recent advances have produced an array of improvements to this approach by decomposing prediction tasks into tailored loss functions that directly target the CATE (e.g., Foster and Syrgkanis (2023); Kennedy (2023); Künzel et al. (2019); Morzywolek et al. (2023); Nie and Wager (2021); van der Laan et al. (2024)). The goal of these approaches is to strip out confounding and baseline effects and then focus on the CATE target parameter. Common in these loss-based approaches are nuisance components that we might leverage for the final CATE goal. In particular, the $R$-learner (Nie and Wager, 2021) leverages Robinson’s transformation (Robins, 2004; Robinson, 1988) to express the CATE function as
@@
One particularly useful ingredient in this loss is the model for the conditional mean in the absence of treatment, $\hat{\mu}_{(0)}(X_{i})$. If features that are predictive of baseline outcomes also matter for treatment benefits, then sharing information between $\hat{\mu}_{(0)}(X_{i})$ and $\hat{\tau}(X_{i})$ might be advantageous in estimating the CATE. One downside of this approach is that it effectively splits the pretraining exercise by treatment assignment, reducing the amount of data available for pretraining from each nuisance parameter. Moreover, pretraining solely on the samples assigned to the control risks overlooking treatment-induced signal in $Y_{i}$, thereby sacrificing a potential safeguard against harmful pretraining, which is automatically built in when pretraining on the mean outcome function in the $R$-Learner.
For these reasons and brevity, we restrict our study to exploring leveraging the $R$-learner to share information between marginalized outcomes $m(X_{i})$ and the CATE. In the supplementary information, we showcase that our logic for pretraining also holds for the $DR$-Learner. For a general discussion and comparison of orthogonal losses for CATE estimation, see, for example, Fisher (2024); Morzywolek et al. (2023).
@@ 2.1 Shared Support between Prognostic and Predictive Factors
A common observation in medicine and other fields is that factors that are prognostic of an outcome are also predictive of treatment response, i.e. there is a overlap in the support of features that are relevant to solve the objective function for the mean outcome function $m(X_{i})$ and those for the CATE function $\tau(X_{i})$. Equation 6 expresses this phenomenon in an easy toy example based on a linear model. The covariate $X_{i1}$ is simply prognostic of the outcome $Y_{i}$ (in the absence of the treatment); $X_{i3}$ is predictive as it interacts with the treatment. Finally, $X_{i2}$ is both prognostic and predictive.
@@
(Prognostic Factor). A prognostic factor is any covariate or combination of covariates that contains information about the likelihood of an outcome, independent of any treatment assignment (Ballman, 2015; National Cancer Institute, 2025b).
@@
Figure 1: Overview of the Pretrained $R$-learner. The figure shows the step-wise procedure of the $R$-learner with the mean outcome function and the CATE. The upper panel shows different pretraining strategies. The lower panels illustrate how these pretraining approaches are integrated into the $R$-learner using the mean outcome function as an example. The active set for the mean outcome model (left panel) is used for pretraining to inform the estimation of the conditional average treatment effect (CATE) model. Here, gender, weight, and biomarker II are not in the support of the mean outcome model (middle panel) and get penalized at $1/\alpha$ when estimating the CATE model (right panel). Optionally, previously known, expert-validated predictive factors are exempt from any penalties. For example, biomarker II is known to be predictive (and thus protected from any penalties a priori), but not in the active set of the mean outcome model. Age, tumor size, and biomarker I are in the support of both models and do not get penalized through pretraining. Icons from Servier Medical Art, licensed under CC BY 4.0, https://creativecommons.org/licenses/by/4.0/ .
@@
The overlap of predictive and prognostic factors also plays a role in psychiatry and cardiology: patients with depression have a higher risk and poorer outcomes for cardiovascular diseases (CVD) (Lichtman et al., 2014). Furthermore, patients who suffer from depression are at higher risk for medication non-adherence, which can decrease the effectiveness of CVD medications such as statins and anti-hypertensive treatments (DiMatteo et al., 2000; Carney and Freedland, 2017). Finally, the phenomenon of shared support is likely to extend to many other disciplines outside medicine. Athey et al. (2025) demonstrates that baseline responses are easier to estimate than treatment effects and leverages this for enhanced causal predictions to target behavioral nudges. In education, for example, school-specific characteristics, such as learning environments, not only predict students’ overall educational performance, but also often modify the effects of educational interventions (Yeager et al., 2019).
@@ 2.2 Side Information from Selection Factors
It is possible to extend our concept of pretraining to the propensity score function in observational settings where treatment assignment is the function of some covariates. Suppose clinicians tend to recommend a treatment to patients in an age group where they believe it is most effective. These patients will have high treatment propensity. If clinicians are generally correct, then the same covariates driving propensity will also predict treatment benefit. Cohen and DeRubeis (2018) discusses the role of prognostic, predictive, and selection factors in the context of depression. Selection factors also play an important role in phase-II and phase-III cancer trials, where investigators choose one (“physician’s choice”) among many treatment options in the control arm (Modi et al., 2022; Ribas et al., 2015).
@@
(Selection Factor). A selection factor is any pre-exposure covariate that contains information about the likelihood of treatment assignment.
Pretraining on selection factors might also be relevant in other areas for research, such as education. There is a large body of research examining the effect of learning interventions or college admission on outcomes (e.g., future income). In this example, selection factors such as socio-economic status, motivation, beliefs about one’s future, and health might be strong effect modifiers (Yeager et al., 2016).
@@ 3 Using the Lasso to Leverage Prognostic Side Information
In order to instantiate an $R$-learner that leverages shared support, we need learners for the mean conditional outcome function and CATE function amenable to this support-sharing task. One tractable solution is to leverage an adaptive lasso for the CATE function, where the differential shrinkage parameters are set through the lasso active set solution for the mean conditional outcome function. For exposition, it is helpful to first consider regularized high-dimensional linear models. Suitable featurizations of the covariate matrix (Hastie et al., 2009) allow us to extend this approach to non-linear settings as revisited in section 4. We start out by assuming a randomized controlled trial setting where the propensity scores are known and omit this estimation task. We start by fitting $m(X_{i})$ with the lasso
@@
where $\lambda_{1}$ is a regularization parameter that minimizes the cross-validated MSE. In addition to the estimated model coefficients $\hat{\beta}$, the lasso yields an active set of features with nonzero coefficients. If some of these coefficients are not part of the active set, then, as argued in Section 2.1, they might be less useful for estimating the CATE. In the next step, we incorporate information from this active set by estimating $\tau(X_{i})$ via an adaptive lasso using (3)
@@
We can employ a similar strategy in the $DR$-learner (5) using its nuisance parameters for pretraining. While pretraining based on the conditional outcome model for the control arm $\mu_{(0)}(X_{i})$ is appealing, it has one disadvantage: depending on the fraction of treated units, we will have less data at our disposal. Moreover, pretraining using $m(X_{i})$ has the potential to pick up signal that stems from $\tau(X_{i})$ given that under Assumption 1 we have that $m(X_{i})=\mu_{(0)}(X_{i})+e(X_{i})\tau(X_{i})$. This provides a potential safeguard against penalizing predictive factors that are not prognostic when there is little to no overlap between the two types of covariates.